# Hessian spectrum at the global minimum of the spherical pure-like mixed $p$-spin glasses

### v2

@@ Abstract
We study the large $N$-dimensional limit of the Hessian spectrum at the global minimum of a subclass of spherical mixed $p$-spin models, extending those results in the regime of topology trivialization.
@@ 1 Introduction
The Hamiltonians of spherical spin glasses are Gaussian smooth functions defined on the sphere in $N$-dimensional Euclidean space. Although the covariance function of the Hamiltonian depends only on the angle between two points, its landscapes are known to be rich and complex. The mean number of critical points (or those with a given index) and topology of its level sets of the Hamiltonian have been rigorously investigated in [ABA13]. On the other hand, the energy landscape near its global maximum was also examined in [Auffinger2018energy]. In this work, we will be interested in the behavior of Hessian spectrum at the global minimum (similarly at the global maximum, by symmetry).
We now introduce the model. Let $S^{N-1}(\sqrt{N})$ be the sphere of radius $\sqrt{N}$ in the $N$-dimensional Euclidean space. The Hamiltonian of the mixed spherical $p$-spin model is given by
@@
where $R\left(\bm{\sigma},\bm{\sigma}^{\prime}\right):=\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}\sigma_{i}^{\prime}$ is the overlap between any two spin configurations $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$. We fix the variance of $H_{N}$ by assuming $\xi(1)=\sum_{p=2}^{\infty}\gamma_{p}^{2}=1.$ If $\xi(x)=x^{p}$ for some $p\geq 3$, the model (1.1) is called the pure $p$-spin model. The assumption (1.2) ensures that the infinite sum in the model (1.1) converges almost surely and the Hamiltonian $H_{N}$ is almost surely smooth and Morse [AT07, Theorem 11.3.1], which guarantees the application of the Kac-Rice formula [AT07, theorem 12.1.1]. Denote the ground state energy in the spherical version of the mixed $p$-spin model by
@@
$F_{N}(\beta)=\frac{1}{\beta N}\log\int_{S^{N-1}(\sqrt{N})}e^{-\beta H_{N}(\bm{\sigma})}\Lambda_{N}(\mathrm{d}\bm{\sigma}),$
where $\Lambda_{N}$ is the normalized surface measure on the sphere $S^{N-1}(\sqrt{N})$. It is well-known that the limiting free energy exists and can be computed through the famous Parisi formula [talagrand2006free, chen2013aizenman]. From here, it can be deduced that the sequence $GS_{N}$ also converges to some value $GS$ almost surely (cf. [AC17, (2) and (3)]), which is usually called limiting ground state energy,
$GS:=\lim_{N\rightarrow\infty}GS_{N},\quad a.s.$
Based on the Crisanti-Sommers representation, a variational formula for the limiting ground state energy $GS$ was also established in [chen], [2016dynamic], which should be seen as the analogue of the Parisi formula at zero temperature.
To understand the landscape of the Hamiltonian $H_{N}(\bm{\sigma})$, Auffinger, Ben Arous, and Černý [ABC13] investigated the mean number of critical points of the spherical pure $p$-spin model. For any Borel set $B\in\mathbb{R}$, denote by $\operatorname{Crt}_{N,k}(B)$ the number of critical points of $H_{N}(\bm{\sigma})$ when the value of $H_{N}(\bm{\sigma})/N$ is restricted to $B$ with index $k$. In this pioneering work, they proved that
@@
where the explicit expression of $\Theta_{p,k}(u)$ is given in [ABC13, (2.16)], which is commonly referred to as the complexity function of the pure $p$-spin model with index $k$. The function $\Theta_{p,k}(u)$ is known to be non-decreasing function. Let $-E_{0}(p)$ be the unique zero of function $\Theta_{p,0}(u)$, and define $E_{\infty}(p)=2\sqrt{\frac{p-1}{p}}$. Moreover, the second moment of $\operatorname{Crt}_{N}((-\infty,u))$ was calculated in [subag2017complexity], which showed that the critical points are concentrated around their means: for any $p\geq 3$ and $u\in\left(-E_{0}(p),-E_{\infty}(p)\right)$,
@@
This result was also obtained earlier in [ABC13, Theorem 2.12] for $p\geq 4$ even. The complexity function was also calculated for the mixed spherical $p$-spin model [ABA13], which states that for any open set $B\in\mathbb{R}$,
@@
$G\left(\xi^{\prime},\xi^{\prime\prime}\right):=\log\frac{\xi^{\prime\prime}(1)}{\xi^{\prime}(1)}-\frac{\left(\xi^{\prime\prime}(1)-\xi^{\prime}(1)\right)\left(\xi^{\prime\prime}(1)-\xi^{\prime}(1)+\xi^{\prime}(1)^{2}\right)}{\xi^{\prime\prime}(1)\xi^{\prime}(1)^{2}}.$
A mixed model with covariance function $\xi(x)$ is called $pure$-$like$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)>0$, $critical$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)=0$, or $full$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)<0$. Denote by $-E_{0}(\xi)$ the smallest zero of $\theta_{0,\xi}(u)$. In Section 4 of their work, they claimed that $\xi(x)$ belonging to the $pure$-$like$ mixture class is equivalent to $E_{0}(\xi)=f_{1}$, where $f_{1}$ can be interpreted as the zero-temperature limit of one-step replica symmetry breaking ($1$-RSB) Parisi functional [ABA13, Propsition 6]. Also through a second moment calculation, Ben Arous, Subag and Zeitouni [BSZ20] proved that $GS=-E_{0}(\xi)$ for the $pure$-$like$ mixture models under Assumption I.
Recently, [Be22] considered the mixed spherical $p$-spin model on the unit sphere $S^{N-1}$ with an external field
@@
where $h\geq 0$ is the strength of the external field and $\mathbf{u}_{N}\in S^{N-1}$ is a deterministic sequence. When $h^{2}>\xi^{\prime\prime}(1)-\xi^{\prime}(1)$, they showed that the Hamiltonian $H_{N}^{h}(\bm{\sigma})$ has trivial geometry, meaning that the only critical points of $H_{N}^{h}$ are a single maximum and a single minimum. It is clear that the property above implies the vanishing of the maximum value of the complexity function. In this setting, they also found the convergence of the largest eigenvalues of the Hessian at global maximum [Be22, Theorem 1.1]. In fact, similar arguments can be used to prove the convergence of the empirical spectrum measures of the Hessian at global maximum, with the help of a large deviation principle (LDP) for the spectrum measures of the GOE matrices [benarous1997large]. Later, [XZ22] also confirmed the trivial topology phenomenon of locally isotropic Gaussian random fields on $\mathbb{R}^{N}$ and studied the convergence of empirical spectrum measures of the Hessian at global minimum provided the value of external field is large, which is previously known in the physics literature [FLD18].
Out of the topology trivialization regime, it may be difficult to find the limiting Hessian spectrum at global minimum, since the maximum value of complexity function is positive, so that with high probability, there are exponentially many critical points. However, it is known that $GS=-E_{0}(p)$ [subag2017complexity, Appendix IV] (see also [chen, Propisition 3]) for the pure $p$-spin model without external fields ($h$=0) and $GS=-E_{0}(\xi)$ [BSZ20, Theorem 1.1] for the $pure$-$like$ mixture models also without external fields (under Assumption I). This means that the smallest zero of the complexity function exactly matches the corresponding $GS$. Together with the analysis of the complexity function, we can study the behavior of the Hessian spectrum at global minimum $\bm{\sigma}^{*}$ in the thermodynamic limit $N\rightarrow\infty$ for these subclass models, which is the main focus of this paper. For more general models, whether this property still holds remains unknown and will require a more detailed investigation in the future. Furthermore, the $full$ mixture models contradict this property above (see [ABA13, Corollary 4.1]), which may pose a significant obstacle to solving this problem.
@@ 1.1 Notations and main results
where $\|f\|_{\infty}$ and $\|f\|_{L}$ denote the $L^{\infty}$ norm and Lipschitz constant of the function $f$, respectively. The empirical spectrum measure of an $N\times N$ matrix $M$ is denoted by $L_{N}(M)=\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}(M)}$, where $\lambda_{i}(M)$ are the eigenvalues of $M$. We also use $\lambda_{\min}(M)$ to denote the smallest eigenvalue of matrix $M$. Let $\sigma_{c,r}$ denote the semicircle measure with center $c$ and radius $r$, which has the following density with respect to Lebesgue measure
@@
In particular, we denote the standard semicircle measure by $\sigma_{\rm sc}=\sigma_{0,2}$. To save space, we sometimes also use the shorthand notation $\xi^{\prime}=\xi^{\prime}(1)$ and $\xi^{\prime\prime}=\xi^{\prime\prime}(1)$.
@@
One can easily verify that the pure $p$-spin model belongs to the $pure$-$like$ mixture class for all $p\geq 3$. However, the particular trivial relation (2.4) exists only in the pure $p$-spin model. Therefore, for the sake of clear presentation, we treat the pure $p$-spin case separately from the $pure$-$like$ but not pure $p$-spin case. For the latter case, we require the following condition as stated in [BSZ20].
@@
Assume that $\xi(x)$ is mixed and $pure$-$like$, that $\frac{d^{2}}{dr^{2}}\Psi_{\xi}^{0}(0)<0$, and that the maximum of $\Psi_{\xi}^{0}(r)$ on the interval $[-1,1]$ is obtained uniquely at $r=0$.
The explicit expression of the function $\Psi_{\xi}^{0}(r)$ is defined in [BSZ20, (1.11)]. Since it is not the key ingredient in this paper, we do not provide it here.
Theorem 1.2 (Pure-like mixture models).
@@
where $y_{0}$ is defined in (3.7).
@@
Recently, one month after we first posted this article on arXiv, Sellek [sellke2024marginal] deduced the limiting Hessian spectrum for even spins (i.e., $\gamma_{p}=0$ for all $p$ odd) via a remarkable construction of the unique minimizer of the Parisi functional of vector $p$-spin model [Zhou2022]. In particular, we emphasize that both our results, Theorems 1.1 and 1.2, include the odd terms. Moreover, the techniques used in our work differ from those in [sellke2024marginal] and should be of independent interest.
We organize the remainder of the paper as follows. We prove the convergence of the empirical spectrum measures and smallest eigenvalues of the Hessian at global minimum for the spherical pure $p$-spin model in Section 2. Section 3 is dedicated to the analysis of spherical $pure$-$like$ mixture models.

### v3

@@ Abstract
We study the large $N$-dimensional limit of the Hessian spectrum at the global minimum of some subclasses of the spherical mixed $p$-spin models. Specifically, we show that its empirical spectral measure converges in probability to a shifted and rescaled semicircle law and does not have outliers. Our method follows the second moment approach developed recently in [BSZ20], from which the ground state energy can be derived for the $pure$-$like$ mixed $p$-spin model. By analyzing the complexity function with given radial derivative and energy, we derive the convergence of the Hessian spectrum from the vanishing mean number of critical points. For the 111-RSB model, the ground state energy was explicitly computed in [huang2023constructive]. Combined with the complexity function of local maxima with given radial derivative obtained in [belius2022complexity], this allows us to obtain the corresponding results more directly. Our result extends those corresponding results in the regime of topology trivialization.
@@ 1 Introduction
The Hamiltonians of spherical spin glasses are Gaussian smooth functions defined on the sphere in $N$-dimensional Euclidean space. Although the covariance function of the Hamiltonian depends only on the angle between two points, its landscapes are known to be rich and complex. The most captivating aspect of spin glass theory is the Parisi formula for the free energy, first predicted by Parisi in 1979 [parisi1979infinite] for the Ising spin models. After decades of progress, Talagrand [talagrand2006parisi] was the first to provide a rigorous mathematical proof of the conjecture aforementioned for the Ising models with even spins, and later extended similar results to the spherical models [talagrand2006free] also with even spins. This formula was further extended to general mixed $p$-spin models by [pan] for the Ising case and by [chen2013aizenman] for the spherical case, respectively. On the other hand, the mean number of critical points (or those with a given index) and topology of its level sets of the Hamiltonian of spherical spin glasses have been rigorously investigated in [ABA13, ABC13] by using the Kac-Rice formula [AT07, theorem 12.1.1]. Since then, the Kac–Rice formula has been applied to study the moments of critical points under various conditions and in different settings; see [subag2017complexity, BSZ20, belius2022complexity, Be22] and the references therein. More recently, Huang and Sellke [huang2023constructive, Lemma 2.8] obtained an ultrametric tree of pure states for the 1-RSB model (see Definition 4.1), each with approximately the ground state energy, through a second moment calculation of the critical points based on the Kac–Rice formula. To study the law of a random field $f_{N}(\bm{\sigma})$ evaluated at a sample from the Gibbs measure associated to the spherical Hamiltonian, Dembo and Subag [dembo2024disordered, Proposition 6.1] derived an upper bound for the complexity of a class of rotationally invariant functions taking values in a special random set via the Kac–Rice formula. In this work, we will be interested in the behavior of Hessian spectrum at the global minimum of some subclasses of the spherical mixed $p$-spin models (similarly at the global maximum, by symmetry).
We now introduce the spherical mixed $p$-spin model. Let $S^{N-1}(\sqrt{N})$ be the sphere of radius $\sqrt{N}$ in the $N$-dimensional Euclidean space. The Hamiltonian of the spherical mixed $p$-spin model is given by
@@
where $R\left(\bm{\sigma},\bm{\sigma}^{\prime}\right):=\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}\sigma_{i}^{\prime}$ is the overlap between any two spin configurations $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$. We fix the variance of $H_{N}$ by assuming $\xi(1)=\sum_{p=2}^{\infty}\gamma_{p}^{2}=1.$ If $\xi(x)=x^{p}$ for some $p\geq 3$, the model (1.1) is called the pure $p$-spin model. The assumption (1.2) ensures that the infinite sum in the model (1.1) converges almost surely and the Hamiltonian $H_{N}$ is almost surely smooth and Morse [AT07, Theorem 11.3.1], which guarantees the application of the Kac-Rice formula. Denote the ground state energy in the spherical version of the mixed $p$-spin model by
@@
$F_{N}(\beta)=\frac{1}{\beta N}\log\int_{S^{N-1}(\sqrt{N})}e^{-\beta H_{N}(\bm{\sigma})}\Lambda_{N}(\mathrm{d}\bm{\sigma}),$ (1.3)
where $\Lambda_{N}$ is the normalized surface measure on the sphere $S^{N-1}(\sqrt{N})$ and the parameter $\beta$ is called the inverse temperature. It is well known that the limiting ground state energy exists almost surely [Lemma 6] [chen], $GS:=\lim_{N\rightarrow\infty}GS_{N}$, which can be deduced by sending the inverse temperature $\beta$ to infinity in the definition of the free energy (1.3). Based on the Crisanti-Sommers representation for free energy, a variational formula for the limiting ground state energy $GS$ was established in [chen, 2016dynamic], which should be regarded as the analogue of the Parisi formula at zero-temperature.
To understand the landscape of the Hamiltonian $H_{N}(\bm{\sigma})$, Auffinger, Ben Arous and Černý [ABC13] firstly investigated the mean number of critical points of the spherical pure $p$-spin model. For any Borel set $B\subset\mathbb{R}$, denote by $\operatorname{Crt}_{N,k}(B)$ the number of critical points of $H_{N}(\bm{\sigma})$ when the value of $H_{N}(\bm{\sigma})/N$ is restricted to $B$ with index $k$. In this pioneering work, they proved that
@@
where the explicit expression of $\Theta_{p,k}(u)$ is given in [ABC13, (2.16)], which is commonly referred to as the complexity function of the pure $p$-spin model with index $k$. Setting $E_{\infty}(p)=2\sqrt{\frac{p-1}{p}}$, the function $\Theta_{p,k}(u)$ is known to be a strictly increasing function on $(-\infty,-E_{\infty}(p))$ and takes the constant value $\frac{1}{2}\log(p-1)-\frac{p-2}{p}>0$ for $u\geq-E_{\infty}(p)$. Let $-E_{0}(p)$ be the unique zero of the function $\Theta_{p,0}(u)$. For any open set $B\subset\mathbb{R}$, denote by $\operatorname{Crt}_{N}(B)$ the number of critical points of $H_{N}(\bm{\sigma})$ when the value of $H_{N}(\bm{\sigma})/N$ is restricted to $B$. Additionally, the second moment of $\operatorname{Crt}_{N}((-\infty,u))$ was calculated in [subag2017complexity], which showed that the number of critical points is concentrated around their means: for any $p\geq 3$ and $u\in\left(-E_{0}(p),-E_{\infty}(p)\right)$,
@@
This result was also obtained earlier in [ABC13, Theorem 2.12] for $p\geq 4$ even. The complexity function was also calculated for the spherical mixed $p$-spin model [ABA13], which states that for any open set $B\subset\mathbb{R}$,
@@
$G\left(\xi^{\prime},\xi^{\prime\prime}\right)=\log\frac{\xi^{\prime\prime}(1)}{\xi^{\prime}(1)}-\frac{\left(\xi^{\prime\prime}(1)-\xi^{\prime}(1)\right)\left(\xi^{\prime\prime}(1)-\xi^{\prime}(1)+\xi^{\prime}(1)^{2}\right)}{\xi^{\prime\prime}(1)\xi^{\prime}(1)^{2}}.$ (1.5)
A mixed model with covariance function $\xi(x)$ is called $pure$-$like$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)>0$, $critical$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)=0$, or $full$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)<0$. Denote by $-E_{0}(\xi)$ the smallest zero of $\theta_{0,\xi}(u)$. In Section 4 of their work, they claimed that $\xi(x)$ belonging to the $pure$-$like$ mixture class is equivalent to $E_{0}(\xi)=f_{1}$, where $f_{1}$ can be interpreted as the zero-temperature limit of 111-RSB Parisi functional [ABA13, Proposition 6]. Also through a second moment calculation, Ben Arous, Subag and Zeitouni [BSZ20] proved that $GS=-E_{0}(\xi)$ for the $pure$-$like$ mixture model under Assumption I.
Recently, [Be22] considered the spherical mixed $p$-spin model on the unit sphere $S^{N-1}$ with an external field
@@
where $h\geq 0$ is the strength of the external field and $\mathbf{u}_{N}\in S^{N-1}$ is a deterministic sequence. When $h^{2}>\xi^{\prime\prime}(1)-\xi^{\prime}(1)$, they showed that the Hamiltonian $H_{N}^{h}(\bm{\sigma})$ has trivial geometry, meaning that the only critical points of $H_{N}^{h}$ are a single maximum and a single minimum. It is clear that the property above implies the vanishing of the maximum value of the complexity function. In this setting, they also found the convergence of the largest eigenvalues of the Hessian at global maximum [Be22, Theorem 1.1]. In fact, similar arguments can be used to prove the convergence of the empirical spectral measures of the Hessian at global maximum, with the help of a large deviation principle (LDP) for the spectral measures of the GOE matrices [benarous1997large]. Later, [XZ22] also confirmed the trivial topology phenomenon for locally isotropic Gaussian random fields defined on $\mathbb{R}^{N}$ and studied the convergence of empirical spectral measures of the Hessian at global minimum provided the value of external field is large, which is previously known in the physics literature [FLD18].
Out of the topology trivialization regime, it may be difficult to find the limiting Hessian spectrum at global minimum, since the maximum value of complexity function can be positive. Subag [subag2021following][Lemma 3] deduced that uniformly for all $\bm{\sigma}\in S^{N-1}(\sqrt{N})$, the Hessian spectrum at $\bm{\sigma}$ can be approximated by a semicircle measure with radius $2\sqrt{\xi^{\prime\prime}(1)}$ and shifted by the rescaled radial derivative $\partial_{r}H_{N}(\bm{\sigma})$, so the key to identifying the Hessian spectrum at global minimum is to determine the radial derivative at global minimum. However, it is known that $GS=-E_{0}(p)$ [subag2017complexity, Appendix IV] (see also [chen, Proposition 3]) for the pure $p$-spin model without external fields ($h$=0) and $GS=-E_{0}(\xi)$ [BSZ20, Theorem 1.1] for the $pure$-$like$ mixture model also without external fields (under Assumption I). This means that the smallest zero of the complexity function exactly matches the corresponding $GS$. By analyzing the complexity function $F(x,y)$ defined in (3), we can prove the convergence of the Hessian spectrum at global minimum $\bm{\sigma}^{*}$ in the thermodynamic limit $N\rightarrow\infty$ for these subclasses models, which is the main focus of this paper. Moreover, for the 1-RSB model, the value of $GS$ can be computed explicitly. Given $x_{*}=-GS$ in the complexity $F(x,y)$ aforementioned, the maximum value of $F(-x_{*},y)$ over variable $y$ is zero, and is obtained uniquely at $-y_{*}$ defined in (4.6), which should represent the corresponding value of the radial derivative at global minimum. Thus, we can also identify the Hessian spectrum at global minimum for this model. For more general models, it is unknown whether this property still holds and will require a more detailed investigation in the future. Instead, the $full$ mixture model contradicts this property (see [ABA13, Corollary 4.1]), which may pose a significant obstacle to solving this problem.
@@ 1.1 Notations and main results
where $\|f\|_{\infty}$ and $\|f\|_{L}$ denote the $L^{\infty}$ norm and Lipschitz constant of the function $f$, respectively. The empirical spectral measure of an $N\times N$ matrix $M$ is denoted by $L_{N}(M)=\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}(M)}$, where $\lambda_{i}(M)$ are the eigenvalues of $M$. We also use $\lambda_{\min}(M)$ and $\lambda_{\max}(M)$ to denote the smallest eigenvalue and the largest eigenvalue of the matrix $M$, respectively. Let $\sigma_{c,r}$ denote the semicircle measure with center $c$ and radius $r$, which has the following density with respect to the Lebesgue measure
@@
In particular, we denote the standard semicircle measure by $\sigma_{\rm sc}=\sigma_{0,2}$. We endow the sphere $S^{N-1}(\sqrt{N})$ with the standard Riemannian structure, induced by the Euclidean Riemannian metric on $\mathbb{R}^{N}$. We fix an orthonormal frame field $E=\left(E_{i}\right)_{i=1}^{N-1}$ on $S^{N-1}(\sqrt{N})$, which means that $\left(E_{i}(\bm{\sigma})\right)_{i=1}^{N-1}$ is a local orthonormal basis of $T_{\bm{\sigma}}S^{N-1}(\sqrt{N})$ for any $\bm{\sigma}\in S^{N-1}(\sqrt{N})$. We then define the spherical gradient and Hessian of $H_{N}(\bm{\sigma})$ as
Furthermore, let $\partial_{r}H_{N}(\bm{\sigma})$ denote the radial derivative of $H_{N}(\bm{\sigma})$ at $\bm{\sigma}\in S^{N-1}(\sqrt{N})$, which can be understood as the directional derivative of $H_{N}(\bm{\sigma})$ in Euclidean space in the direction $\bm{\sigma}$. For the relationship between $H_{N}(\bm{\sigma})$ and its derivatives, we refer the reader to [Be22, Lemma 3.2]. To save space, we sometimes also use the shorthand notation $\xi^{\prime}=\xi^{\prime}(1)$ and $\xi^{\prime\prime}=\xi^{\prime\prime}(1)$.
@@
$\lim_{N\rightarrow\infty}\lambda_{\max}\left(\nabla^{2}_{\mathrm{sp}}H_{N}\left(\bm{\sigma}^{*}\right)\right)=c_{p}+r_{p},$ (1.9)
@@
One can easily verify that the pure $p$-spin model belongs to the $pure$-$like$ mixture class for any $p\geq 3$. However, the specific trivial relation (2.4) exists only in the pure $p$-spin model. Therefore, for the sake of clear presentation, we treat the pure $p$-spin case separately from the $pure$-$like$ but not pure $p$-spin case. For the latter case, we require the following condition as stated in [BSZ20].
@@
Assume that $\xi(x)$ is mixed and $pure$-$like$, that $\frac{d^{2}}{dr^{2}}\Psi_{\xi}^{0}(0)<0$, and that the maximum of $\Psi_{\xi}^{0}(r)$ on the interval [1,1][-1,1][ - 1 , 1 ] is obtained uniquely at $r=0$.
The explicit expression of the function $\Psi_{\xi}^{0}(r)$ is defined in [BSZ20, (1.11)]. Since it is not the key ingredient in this paper, we do not provide it here. This assumption ensures that zero overlap is the unique maximizer of the pair complexity function at levels $-E_{0}$ and $-y_{0}$ as defined in (3.7). Additionally, it is used to justify the application of the second moment methods, which can be employed to identify the ground state energy. [BSZ20, Theorem 3] deduced that under this assumption, the model lies in 111-RSB phase at zero-temperature, but this assumption does not cover the entire zero-temperature 111-RSB phase [huang2023constructive, Lemma 3.5].
Theorem 1.2 ($Pure$-$like$ mixture model).
@@
$\lim_{N\rightarrow\infty}\lambda_{\max}\left(\nabla^{2}_{\mathrm{sp}}H_{N}\left(\bm{\sigma}^{*}\right)\right)=y_{0}+2\sqrt{\xi^{\prime\prime}}$ (1.12)
where $y_{0}>2\sqrt{\xi^{\prime\prime}}$ is defined in (3.7).
@@
Recently, one month after we first posted this article on arXiv, Sellke [sellke2024marginal] deduced the limiting Hessian spectrum for even spins (i.e., $\gamma_{p}=0$ for all $p$ odd) via a remarkable construction of the unique minimizer of the Parisi functional of vector $p$-spin model [Zhou2022]. In particular, we emphasize that all of our results, Theorems 1.1, 1.2 and 1.3 can include the odd terms. Furthermore, the techniques used in our work differ from those in [sellke2024marginal] and should be of independent interest.
Recall the definition of 111-RSB model in Definition 4.1 and the inequality $y_{*}\geq 2\sqrt{\xi^{\prime\prime}}$ given in (4.6). For this model, we have the following result.
Theorem 1.3 (111-RSB model).
Assume $\xi(x)$ is 111-RSB and $y_{*}>2\sqrt{\xi^{\prime\prime}}$. Let $\bm{\sigma}^{*}$ denote the global minimum of $H_{N}(\bm{\sigma})$. Then we have convergence in probability
$\lim_{N\to\infty}d(L_{N-1}(\nabla^{2}_{\mathrm{sp}}H_{N}(\bm{\sigma}^{*})),\sigma_{y_{*},2\sqrt{\xi^{\prime\prime}}})=0,$ (1.13)
$\lim_{N\rightarrow\infty}\lambda_{\min}\left(\nabla^{2}_{\mathrm{sp}}H_{N}\left(\bm{\sigma}^{*}\right)\right)=y_{*}-2\sqrt{\xi^{\prime\prime}},$ (1.14)
$\lim_{N\rightarrow\infty}\lambda_{\max}\left(\nabla^{2}_{\mathrm{sp}}H_{N}\left(\bm{\sigma}^{*}\right)\right)=y_{*}+2\sqrt{\xi^{\prime\prime}}.$ (1.15)
Remark 1.2.
Here we slightly strengthen the inequality to $y_{*}>2\sqrt{\xi^{\prime\prime}}$, as this is required to apply [Be22, Lemma 6.4] in order to obtain an upper bound for the second moment of the absolute value of the determinant of a shifted GOE matrix. The lemma requires that the shift lies slightly outside the support of the standard semicircle law $\sigma_{0,2}$. This condition $y_{*}>2\sqrt{\xi^{\prime\prime}}$ is equivalent to $y>\sqrt{\xi^{\prime\prime}}$, as defined in (2.14).
We organize the remainder of this paper as follows. We prove the convergence of the empirical spectral measures, smallest eigenvalues and largest eigenvalues of the Hessian at global minimum for the spherical pure $p$-spin model in Section 2. Section 3 and Section 4 are dedicated to the analysis of spherical $pure$-$like$ mixture model and 111-RSB model, respectively.